# Supervariable Approach

### 16 Feb 2024 v2

@@ 1 Introduction
For the canonical quantization of constrained systems, two different approaches involving the Lagrangian and Hamiltonian of the system have been adapted. The formulation of canonical quantization using Hamiltonian of the system has been established by Dirac in his seminal papers [26, 27, 28]. In Dirac’s canonical quantization of the reparametrization invariant theories, the wavefunction representing the physical state of the system must be annihilated by the operator form of Hamiltonian constraint and, thus, consequently leads to the well-known Wheeler-De-Witt equation [29]. In order to covariantly quantize a system, having local gauge symmetry, the removal of redundant degrees of freedom requires to fix the gauge completely. The latter breaks the inherent local gauge symmetry of the system. For the restoration of symmetry, the phase space must be extend. One of the ways to extend the phase space, by adding extra degrees of freedom, is the Becchi–Route–Stora–Tyutin (BRST) formalism [30, 31, 32, 33]. In this formulation, the gauge-fixing and Faddeev-Popov (anti-)ghost terms are incorporated into the action to ensure the global supersymmetric type BRST invariance. In the BRST quantization, the addition of Faddeev-Popov (anti-)ghost term does not change the physical content of the effective theory. It is well established that for a given a local gauge symmetry, in addition to global BRST symmetry, we also have global anti-BRST symmetry. Both BRST and anti-BRST symmetry transformations are nilpotent of order two and absolutely anticommuting in nature [34, 35]. Subsequently, against the backdrop of BRST invariance, other approaches such as Batalin–Fradkin–Vilkovisky (BFV) [36, 37] and Batalin–Vilkovisky (BV) have been formulated [38].
@@
One of the elegant and geometrically intuitive methods for obtaining the (anti-)BRST symmetry transformations corresponding to a local gauge symmetry is the Bonora-Tonin (BT) superfield formalism [39, 40]. In the BT superfield approach, the (anti-)BRST transformations found their geometrical interpretation in terms of the Grassmannian translational generators. In this geometrical superfield formulation, one extends a general $D$-dimensional Minkowskian spacetime manifold parametrized by the bosonic coordinates $x^{\mu}$ (with $\mu=0,1,2,...,(D-1)$) to the $(D,2)$-dimensional supermanifold ($Z^{M}=(x^{\mu},\theta,\bar{\theta}))$ parametrized by two additional Grassmannian variables $\theta$ and $\bar{\theta}$. These Grassmannian variables, by construction, are nilpotent of order two ($\theta^{2}=\bar{\theta}^{2}=0$) as well as anticommuting ($\theta\bar{\theta}+\bar{\theta}\theta=0$). In addition, for example, the ordinary one-form connection $A^{(1)}=dx^{\mu}A_{\mu}$ and exterior derivative $d=dx^{\mu}\partial_{\mu}$ are generalised to super-one-form connection ${\cal A}^{(1)}=dZ^{M}{\cal A}_{M}=dx^{\mu}{\cal A}_{\mu}+d\theta\,\bar{\cal C}+\bar{d}\theta\,{\cal C}$ and super-exterior derivative ${\widetilde{d}}=dx^{\mu}\,\partial_{\mu}+d\theta\,\partial_{\theta}+d\bar{\theta}\,\partial_{\bar{\theta}}$ onto $(D,2)$-dimensional supermanifold, respectively. One of the important features of the superfield approach BRST formalism is the natural emergence of Curci-Ferrari (CF) condition [34]. This CF condition is required for the proof of absolute anticommutativity of (anti-)BRST transformations as well as (anti-)BRST invariance of the gauge-fixed coupled Lagrangian densities. In addition, CF condition also respects the BRST and anti-BRST transformations. In recent works [41, 42], the superfield formalism has been applied to a class of local reparametrization (or diffeomorphism) invariant field-theoretic models and derived the (anti-)BRST symmetry transformations where the (super-)diffeomorphism invariance play a decisive role.
@@
In our present endeavour, we consider a generally covariant system of gravity in ADM formulation [17]. One such system is the FLRW model of the homogeneous and isotropic Universe [45, 46]. From the Lagrangian point of view, the FLRW model respects the time-reparametrization (i.e. $(0+1)$-dimensional diffeomorphism) symmetry transformation. The path integral quantization and the derivation of Wheeler-De-Witt equation for such time-reparametrization invariant models were carried out by Halliwell [43, 44]. The quantum theory of FLRW model has been studied within the context of BV formalism, where the lapse function and scale factor are incorporated in the Lagrangian [45, 46]. In FLRW model, the lapse function and scale factor are known as ADM variables. Earlier, a few inconsistencies between Dirac and ADM approaches were pointed out [47] but in the context of BRST quantization these have been resolved by extending the phase space where the physical parameters and gauge degrees of freedom are taken into account at equal footing [45, 46]. Furthermore, a suitable scalar product has been defined which ensures that the super-Hamiltonian for FLRW model vanishes. In addition, a connection between the quantization of FLRW model by using the Lagrangian and Dirac’s formulations of constrained dynamics has been established [48].
@@
The layout of our present endeavour is organized as follows. In Section 2, we discuss very briefly about the FLRW model and associated gauge and reparametrization symmetries. Our Section 3 deals with the derivation of off-shell nilpotent and anticommuting (anti-)BRST symmetry transformations corresponding to a local time-reparametrization transformation in the context of supervarable approach to BRST formalism. Section 4 is devoted to the discussion on the nilpotency and anticommutativity properties of the (anti-) BRST transformation as well as the validity of the CF-type condition. The (anti-)BRST invariance of the CF-type of condition is shown in this section, too. By exploiting the basic tenets of BRST formalism, we derive the coupled (but equivalent) Lagrangians in Section 5. In Section 6, we conclude the present work by summarizing our results and discuss the implication of our findings for the future investigation. Finally, in Appendix Appendix A: On the derivarion of CF-type of condition, we provide an explicit derivation of the CF-type of restriction.
@@ 2 Gauge and Reparmetrization Symmetries of FLRW Model
where the lapse function $N(t)$ and scale factor $a(t)$ explicitly depend on the time-evolution parameter ($t$). It is to be noted that the parameter $\kappa=1,0,-1$ represents the closed, flat and open Universe, respectively. In ADM formulation, the Lagrangian corresponding to FLRW metric in $(0+1)$-dimensional (1D) spacetime can be written as [45, 46]
@@
where, $\Omega_{1}$ denotes the primary constraint on the theory and $\pi_{N}=\partial L/\partial\dot{N}=0$ defines the canonical momentum corresponding to the generalized variable $N$. Using the Legendre transformations, the canonical Hamiltonian $H_{c}$ of the system reads
@@
As a consequence of above, the present model is a gauge-theoretic model endowed with first-class constraints. Moreover, these first-class constraints are the generators of local gauge symmetry transformations. For the present FLRW model, the gauge symmetry transformations generator is given as [50]
@@
yield the Lagrangian (6) quasi-invariant. In other words, Lagrangian transforms to a total time derivative as one can check
@@ 3 Derivation of proper (anti-)BRST transformations: Supervariable approach
where, $\Phi(t,\theta,\bar{\theta})={\cal A}(t,\theta,\bar{\theta}),\;\Pi_{\cal A}(t,\theta,\bar{\theta})$. In other words, the coefficients of $\bar{\theta}$ and $\theta$ in the super-expansions of the supervariables (cf. Eq. (16)) are nothing but the BRST $(s_{b})$ and anti-BRST $(s_{ab})$ transformations, respectively. Therefore, in the operator form, we have the mappings between the (anti-)BRST transformations $(s_{(a)b})$ and Grassmannian translational generators $\big{(}\frac{\partial}{\partial\theta},\;\frac{\partial}{\partial\bar{\theta}}\big{)}$, namely; $s_{b}\leftrightarrow\frac{\partial}{\partial\bar{\theta}}\Big{|}_{\theta=0}$ and $s_{ab}\leftrightarrow\frac{\partial}{\partial\theta}\Big{|}_{\bar{\theta}=0}$.
@@
where, we have purposely chosen $B=-(h+\dot{\bar{C}}\,C)$. We shall see later on, that $B$ will play the role of Nakanishi-Lautrup auxiliary variable of the BRST-invariant theory. It is to be noted that the terms in square brackets represent the (anti-)BRST transformations for the generic variable $\phi=a,\,\pi_{a}$.
@@
For the derivation of proper (anti-)BRST transformations of lapse function $N(t)$, we recall the property of one-dimensional differomphism i.e. $N(t)$ transforms like a density function under time-reparametrization i.e. $N^{\prime}(t^{\prime})dt^{\prime}=N(t)\,dt$. In the language of differential form, the density function $N(t)\,dt=N^{(1)}$ defines a one-form on $(0+1)$-dimensional manifold. Thus, the corresponding super-one-form ${\cal N}^{(1)}={\cal N}_{M}(T)\,dT^{M}$ onto $(1,2)D$ supemanifold is given as
@@
where, ${\widetilde{d}}=dt\,\partial/\partial t+d\theta\,\partial/\partial\theta+d\bar{\theta}\,\partial/\partial{\bar{\theta}}$ defines the super-exterior derivative onto $(1,2)D$ supemanifold corresponding to an ordinary exterior derivative $d=dt\,\partial/\partial t$ defined onto $(0+1)D$ manifold. In the above, the quantity ${\widetilde{d}}\,{\widetilde{T}}^{M}$ can be written in matrix form as
@@
where, for the sake of brevity, we have chosen $B=-(h+\dot{\bar{C}}\,C)$ and $\bar{B}=(h+\bar{C}\,\dot{C})$ which play the role of Nakanishi-Lautrup type auxiliary variables. After making these choices for the Nakanishi-Lautrup type auxiliary variables, the super-expansions of the supervariables will take the simplest form (cf. Eqs. (30) and (32). As a result, we obtain the standard and desired (anti-)BRST transformations (see, Section 4 below).
@@
Before we wrap up this subsection, a couple of points is in order. First, we point out that the algebraic relationships in the last line of Eq. (29) yield the celebrated CF-type conditionFor the explicit derivation of CF-type condition, see Appendix Appendix A: On the derivarion of CF-type of condition.
@@
This CF-type condition, as we shall see later, play a crucial role in the absolute anticommutativity of the BRST and anti-BRST transformations (see, Eq. (37) below). And, it will also play an important in the derivation of coupled Lagrangians which respect both BRST and anti-BRST transformations (see, Section 5 below). Second, the (anti-)BRST transformations of the Nakanishi-Lautrup type auxiliary variables $B$ and ${\bar{B}}$ have not yet been determined because we do not know the expressions of supervariables corresponding to the Nakanishi-Laurtup type auxiliary variables. These variables appeared in the super-expansions of $(\bar{\cal C}){\cal C}$ as the secondary variables (cf. Eq. (32)) and they are also related to the secondary variable $h$ defined in the super-expansion of ${\cal F}(T)={\cal F}(t,\theta,\bar{\theta})$ (cf. Eq. (14)). Although, the super-expansions for $B$ and $\bar{B}$ can be written, if the (anti-)BRST transformations of $B$ and $\bar{B}$ are known. We shall determine the (anti-)BRST transformations for $B$ and $\bar{B}$ using the key properties (i.e. nilpotency and absolute anticommutativity properties) of (anti-)BRST transformations in our upcoming Section 4.
@@
modulo total time derivative. Here, the subscripts $(a),\,(a)b$ on the Lagrangian denote the BRST and anti-BRST invariant Lagrangians, respectively. The equations of motion derived from the above Lagrangians w.r.t. the auxiliary variables $B$ and $\bar{B}$ enforce the differential gauge-fixing condition $\dot{N}-\dot{\chi}(a)=0$ (cf. Eq. (44)). It is straightforward to check that both $L_{(b)}$ and $L_{(ab)}$ are coupled Lagrangians because the Nakanishi-Laurtup type auxiliary variables $B$ and $\bar{B}$ are related through CF-tpye condition (33). Again, both Lagrangians differ by an algebraic relation, as one can readily check that
@@
$s_{ab}\,L_{(ab)}$ $=$ $\dfrac{d}{dt}\,\Big{[}\Big{(}\dot{a}\,\pi_{a}+\dfrac{N}{2}\Big{(}\dfrac{\pi^{2}_{a}}{a}+\kappa a\Big{)}\Big{)}\bar{C}-\bar{B}\,\big{(}\dot{N}-\dot{\chi}(a)\big{)}\bar{C}-\bar{B}\,N\dot{\bar{C}}\Big{]},$ (49)
manifesting the corresponding actions remain invariant for the physically well-defined variables which vanish-off rapidly at $t=\pm\infty$. As mentioned above, both Lagrangian are equivalent on the constraint hypersurface defined by Eq. (33), the Lagrangians $L_{(b)}$ and $L_{(ab)}$ must respect the anti-BRST and BRST transformations, respectively on the constraint hypersurface (33). The application of $s_{ab}$ and $s_{b}$ on $L_{(b)}$ and $L_{(ab)}$, respectively yield
@@
where, ${\cal A}^{-1}(t,\theta,\bar{\theta})$ is the inverse of ${\cal A}(t,\theta,\bar{\theta})$. It is evident form Eq. (53), the (anti-)BRST invariance of first-order Lagrangian can be expressed in terms of Grassmannian translational generators $\partial/\partial\theta,\;\partial/\partial{\bar{\theta}}$ and supervariables as
@@
where, the supervariable ${\cal X}({\cal A}(t,\theta,\bar{\theta}))$ defined onto $(1,2)$-dimensional supermanifold is the generalization of an ordinary function $\chi(a)$. The former is given by the following expression

### 05 Mar 2024 v3

@@ 1 Introduction
For the canonical quantization of constrained systems, two different approaches involving the Lagrangian and Hamiltonian of the system have been adapted. The formulation of canonical quantization using Hamiltonian of the system has been established by Dirac in his seminal papers [26, 27, 28]. In Dirac’s canonical quantization of the reparametrization invariant theories, the wavefunction representing the physical state of the system must be annihilated by the operator form of Hamiltonian constraint and, thus, consequently leads to the well-known Wheeler-DeWitt equation [29]. In order to covariantly quantize a system, having local gauge symmetry, the removal of redundant degrees of freedom requires to fix the gauge completely. The latter breaks the inherent local gauge symmetry of the system. For the restoration of symmetry, the phase space must be extend. One of the ways to extend the phase space, by adding extra degrees of freedom, is the Becchi–Rouet–Stora–Tyutin (BRST) formalism [30, 31, 32, 33]. In this formulation, the gauge-fixing and Faddeev-Popov (anti-)ghost terms are incorporated into the action to ensure the global supersymmetric type BRST invariance. In the BRST quantization, the addition of Faddeev-Popov (anti-)ghost term does not change the physical content of the effective theory. It is well established that for a given local gauge symmetry, in addition to global BRST symmetry, we also have global anti-BRST symmetry. Both BRST and anti-BRST symmetry transformations are nilpotent of order two and absolutely anticommuting in nature [34, 35]. Subsequently, against the backdrop of BRST invariance, other approaches such as Batalin–Fradkin–Vilkovisky (BFV) [36, 37] and Batalin–Vilkovisky (BV) have been formulated [38].
@@
One of the elegant and geometrically intuitive methods for obtaining the (anti-)BRST symmetry transformations corresponding to a local gauge symmetry is the Bonora-Tonin (BT) superfield formalism [39, 40]. In the BT superfield approach, the (anti-)BRST transformations found their geometrical interpretation in terms of the Grassmannian translational generators. In this geometrical superfield formulation, one extends a general $D$-dimensional Minkowskian spacetime manifold parametrized by the bosonic coordinates $x^{\mu}$ (with $\mu=0,1,2,...,(D-1)$) to the $(D,2)$-dimensional supermanifold ($Z^{M}=(x^{\mu},\theta,\bar{\theta}))$ parametrized by two additional Grassmannian variables $\theta$ and $\bar{\theta}$. These Grassmannian variables, by construction, are nilpotent of order two ($\theta^{2}=\bar{\theta}^{2}=0$) as well as anticommuting ($\theta\bar{\theta}+\bar{\theta}\theta=0$). In addition, for example, the ordinary one-form connection $A^{(1)}=dx^{\mu}A_{\mu}$ and exterior derivative $d=dx^{\mu}\partial_{\mu}$ are generalised to super-one-form connection ${\cal A}^{(1)}=dZ^{M}{\cal A}_{M}=dx^{\mu}{\cal A}_{\mu}+d\theta\,\bar{\cal C}+\bar{d}\theta\,{\cal C}$ and super-exterior derivative ${\widetilde{d}}=dx^{\mu}\,\partial_{\mu}+d\theta\,\partial_{\theta}+d\bar{\theta}\,\partial_{\bar{\theta}}$ onto $(D,2)$-dimensional supermanifold, respectively. One of the important features of the superfield approach to BRST formalism is the natural emergence of Curci-Ferrari (CF) condition [34]. This CF condition is required for the proof of absolute anticommutativity of (anti-)BRST transformations as well as (anti-)BRST invariance of the gauge-fixed coupled Lagrangian densities. In addition, CF condition also respects the BRST and anti-BRST transformations. In recent works [41, 42], the superfield formalism has been applied to a class of local reparametrization (or diffeomorphism) invariant field-theoretic models and derived the (anti-)BRST symmetry transformations where the (super-)diffeomorphism invariance play a decisive role.
@@
In our present endeavour, we consider a generally covariant system of gravity in ADM formulation [17]. One such system is the FLRW model of the homogeneous and isotropic Universe [45, 46]. From the Lagrangian point of view, the FLRW model respects the time-reparametrization (i.e. $(0+1)$-dimensional diffeomorphism) symmetry transformation. The path integral quantization and the derivation of Wheeler-DeWitt equation for such time-reparametrization invariant models were carried out by Halliwell [43, 44]. The quantum theory of FLRW model has been studied within the context of BV formalism, where the lapse function and scale factor are incorporated in the Lagrangian [45, 46]. In FLRW model, the lapse function and scale factor are known as ADM variables. Earlier, a few inconsistencies between Dirac and ADM approaches were pointed out [47] but in the context of BRST quantization these have been resolved by extending the phase space where the physical parameters and gauge degrees of freedom are taken into account at equal footing [45, 46]. Furthermore, a suitable scalar product has been defined which ensures that the super-Hamiltonian for FLRW model vanishes. In addition, a connection between the quantization of FLRW model by using the Lagrangian and Dirac’s formulations of constrained dynamics has been established [48].
@@
The layout of our present endeavour is organized as follows. In Section 2, we discuss very briefly about the FLRW model and associated gauge and reparametrization symmetries. Our Section 3 deals with the derivation of off-shell nilpotent and anticommuting (anti-)BRST symmetry transformations corresponding to a local time-reparametrization transformation in the context of supervarable approach to BRST formalism. Section 4 is devoted to the discussion on the nilpotency and anticommutativity properties of the (anti-) BRST transformation as well as the validity of the CF-type condition. The (anti-)BRST invariance of the CF-type of condition is shown in this section, too. By exploiting the basic tenets of BRST formalism, we derive the coupled (but equivalent) Lagrangians in Section 5. In Section 6, we conclude the present work by summarizing our results and discuss the implication of our findings for the future investigation. Finally, in Appendix A, we provide an explicit derivation of the CF-type of restriction.
@@ 2 Gauge and Reparmetrization Symmetries of FLRW Model
where the lapse function $N(t)$ and scale factor $a(t)$ explicitly depend on the time-evolution parameter $t$. It is to be noted that the parameter $\kappa=1,0,-1$ represents the closed, flat and open Universe, respectively. In ADM formulation, the Lagrangian corresponding to FLRW metric in $(0+1)$-dimensional (1D) spacetime can be written as [45, 46]
@@
where $\Omega_{1}$ denotes the primary constraint on the theory and $\pi_{N}=\partial L/\partial\dot{N}=0$ defines the canonical momentum corresponding to the generalized variable $N$. Using the Legendre transformations, the canonical Hamiltonian $H_{c}$ of the system reads
@@
As a consequence of the above, the present model is a gauge-theoretic model endowed with first-class constraints. Moreover, these first-class constraints are the generators of the local gauge symmetry transformations. For the present FLRW model, the gauge symmetry transformations generator is given as [50]
@@
which leave the Lagrangian (6) quasi-invariant. In other words, Lagrangian transforms to a total time derivative as one can check
@@ 3 Derivation of proper (anti-)BRST transformations: Supervariable approach
where $\Phi(t,\theta,\bar{\theta})={\cal A}(t,\theta,\bar{\theta}),\;\Pi_{\cal A}(t,\theta,\bar{\theta})$. In other words, the coefficients of $\bar{\theta}$ and $\theta$ in the super-expansions of the supervariables (cf. Eq. (16)) are nothing but the BRST $(s_{b})$ and anti-BRST $(s_{ab})$ transformations, respectively. Therefore, in the operator form, we have the mappings between the (anti-)BRST transformations $(s_{(a)b})$ and Grassmannian translational generators $\big{(}\frac{\partial}{\partial\theta},\;\frac{\partial}{\partial\bar{\theta}}\big{)}$, namely; $s_{b}\leftrightarrow\frac{\partial}{\partial\bar{\theta}}\Big{|}_{\theta=0}$ and $s_{ab}\leftrightarrow\frac{\partial}{\partial\theta}\Big{|}_{\bar{\theta}=0}$.
@@
where we have purposely chosen $B=-(h+\dot{\bar{C}}\,C)$. We shall see later on, that $B$ will play the role of Nakanishi-Lautrup auxiliary variable of the BRST-invariant theory. It is to be noted that the terms in square brackets represent the (anti-)BRST transformations for the generic variable $\phi=a,\,\pi_{a}$.
@@
For the derivation of proper (anti-)BRST transformations of the lapse function $N(t)$, we recall the property of one-dimensional differomphism i.e. $N(t)$ transforms like a density function under time-reparametrization (i.e. $N^{\prime}(t^{\prime})dt^{\prime}=N(t)\,dt$). In the language of differential form, the density function $N(t)\,dt=N^{(1)}$ defines a one-form onto $(0+1)$-dimensional manifold. Thus, the corresponding super-one-form ${\cal N}^{(1)}={\cal N}_{M}(T)\,dT^{M}$ onto $(1,2)D$ supemanifold is given as
@@
where ${\widetilde{d}}=dt\,\partial/\partial t+d\theta\,\partial/\partial\theta+d\bar{\theta}\,\partial/\partial{\bar{\theta}}$ defines the super-exterior derivative onto $(1,2)D$ supemanifold corresponding to an ordinary exterior derivative $d=dt\,\partial/\partial t$ defined onto $(0+1)D$ manifold. In the above, the quantity ${\widetilde{d}}\,{\widetilde{T}}^{M}$ can be written in matrix form as
@@
where, for the sake of brevity, we have chosen $B=-(h+\dot{\bar{C}}\,C)$ and $\bar{B}=(h+\bar{C}\,\dot{C})$ which play the role of Nakanishi-Lautrup type auxiliary variables. After making these choices for the Nakanishi-Lautrup type auxiliary variables, the super-expansions of the supervariables will take the simplest form (cf. Eqs. (30) and (32)). As a result, we obtain the standard and desired (anti-)BRST transformations (see, Section 4 below).
@@
Before we wrap up this subsection, a couple of points are in order. First, we point out that the algebraic relationships in the last line of Eq. (29) yield the celebrated CF-type conditionFor the explicit derivation of CF-type condition, see Appendix A.
@@
This CF-type condition, as we shall see later, plays a crucial role in the absolute anticommutativity of the BRST and anti-BRST transformations (see, Eq. (37) below). And, it will also play an important in the derivation of coupled Lagrangians which respect both BRST and anti-BRST transformations (see, Section 5 below). Second, the (anti-)BRST transformations of the Nakanishi-Lautrup type auxiliary variables $B$ and ${\bar{B}}$ have not yet been determined because we do not know the expressions of supervariables corresponding to the Nakanishi-Laurtup type auxiliary variables. These variables appeared in the super-expansions of $(\bar{\cal C}){\cal C}$ as the secondary variables (cf. Eq. (32)) and they are also related to the secondary variable $h$ defined in the super-expansion of ${\cal F}(T)={\cal F}(t,\theta,\bar{\theta})$ (cf. Eq. (14)). Although, the super-expansions for $B$ and $\bar{B}$ can be written, if the (anti-)BRST transformations of $B$ and $\bar{B}$ are known. We shall determine the (anti-)BRST transformations for $B$ and $\bar{B}$ using the key properties (i.e. nilpotency and absolute anticommutativity properties) of (anti-)BRST transformations in our upcoming Section 4.
@@
modulo total time derivative. Here, the subscripts $(a),\,(ab)$ on the Lagrangians denote the BRST and anti-BRST invariant Lagrangians, respectively. The equations of motion derived from the above Lagrangians w.r.t. the auxiliary variables $B$ and $\bar{B}$ enforce the differential gauge-fixing condition $\dot{N}-\dot{\chi}(a)=0$ (cf. Eq. (44)). It is straightforward to check that both $L_{(b)}$ and $L_{(ab)}$ are coupled Lagrangians because the Nakanishi-Laurtup type auxiliary variables $B$ and $\bar{B}$ are related through CF-tpye condition (33). Again, both Lagrangians differ by an algebraic relation, as one can readily check that
@@
$s_{ab}\,L_{(ab)}$ $=$ $\dfrac{d}{dt}\,\Big{[}\Big{(}\dot{a}\,\pi_{a}+\dfrac{N}{2}\Big{(}\dfrac{\pi^{2}_{a}}{a}+\kappa a\Big{)}\Big{)}\bar{C}-\bar{B}\,\big{(}\dot{N}-\dot{\chi}(a)\big{)}\bar{C}-\bar{B}\,N\dot{\bar{C}}\Big{]}.$ (49)
Thus, the corresponding actions remain invariant for the physically well-defined variables which vanish-off rapidly at $t=\pm\infty$. As mentioned above, both Lagrangian are equivalent on the constraint hypersurface defined by Eq. (33), the Lagrangians $L_{(b)}$ and $L_{(ab)}$ must respect the anti-BRST and BRST transformations, respectively on the constraint hypersurface (33). The application of $s_{ab}$ and $s_{b}$ on $L_{(b)}$ and $L_{(ab)}$, respectively yield
@@
where ${\cal A}^{-1}(t,\theta,\bar{\theta})$ is the inverse of ${\cal A}(t,\theta,\bar{\theta})$. It is evident form Eq. (53), the (anti-)BRST invariance of first-order Lagrangian can be expressed in terms of Grassmannian translational generators $\partial/\partial\theta,\;\partial/\partial{\bar{\theta}}$ and supervariables as
@@
where the supervariable ${\cal X}({\cal A}(t,\theta,\bar{\theta}))$ defined onto $(1,2)$-dimensional supermanifold is the generalization of an ordinary function $\chi(a)$. The former is given by the following expression